]the singles court in tennis is 4 yards longer than it is wide. if its perimeter is 44 yards, then what are the length and width?]

Found 2 solutions by TimothyLamb, tutoringisfun:
Answer by TimothyLamb(4379)

L = w + 4
2L + 2w = 44
---
2L + 2w = 44
2(w + 4) + 2w = 44
2w + 8 + 2w = 44
4w = 36
---
w = 9 yards
L = 13 yards
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php

Answer by tutoringisfun(17)

Good afternoon student. I hope this is helpful to you. Thank you for choosing Algebra.com.

1. Write down pertinent information from the problem.
P=44
L=w+4
2. Although it is not given in the problem, one needs to use the following formula for determining the perimeter of a rectangle
P=2L+2W
3. Plug in the values given in the problem and solve:
44=2(w+4)+2W

distribute the 2 to get rid of the parentheses:
44=2W+8+2W
combine like terms:
44=4W+8
subtract 8 from each side because we need to get w by itself
44-8=4W+8-8
on the right, the 8's cancel out leaving:
36=4W
Divide both sides by 4
36/4=4W/4
Solve
9=W
4. Solve the following equation for L
L=W+4
substitute 9 for W and add
L=9+4
13=L
5. to check your answer plug the numbers for L,W into the formula for perimeter from above
P=2L+2W
44=2(13)+2(9)
44=26+18
44=44
I hope this was helpful to you. If you desire PAID assistance, you can contact me via the following websites
www.instaedu.com
www.eduboard.com
or via my facebook page as mentioned in my profile. Good Luck in school.